OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y=-1 x-3y=11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

A

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

2-3=-1 -1=-1 2-3(-3)=11 2+9=11 (note that negative times negative equals positive) 11=11

OpenStudy (anonymous):

Just substitute the values remember an ordered pair is (x,y)

OpenStudy (anonymous):

omg lol I forgot all about negative times negative is positive

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Complete the explanation that describes a linear function for which the least value in the range does not occur at the least value of the domain (a function for which the least value in the range is not the initial value). A line segment that has a __________________ (positive, negative) slope has an initial value that is the greatest value in the range.